What is the average kinetic energy of a molecule of oxygen at a temperature of 330 K?
Answer #1

Here ,

T = 330 K

for the average kinetic energy of a molecule

K = 5 * 0.50 * k * T

K = 5 * 0.50 * 1.38 *10^-23 * 330

K = 1.139 *10^-20 J

the average kinetic energy of molecule of oxygen is 1.139 *10-20 J